# EquiContact: A Hierarchical SE(3) Vision-to-Force Equivariant Policy for Spatially Generalizable Contact-rich Tasks

### v1

@@ Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant ACT, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on PiH tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
@@ I Introduction
However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust policies. As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations.
@@
[…]
In this paper, we propose EquiContact, a hierarchical SE(3) vision-to-force equivariant policy for spatially generalizable contact-rich tasks. Our proposed method consists of two main components. The first is a high-level planner consisting of a Diff-EDF model [14], which uses global point cloud data to provide a local reference frame for placing the peg relative to the hole. The second, low-level compliant visuomotor policy is a variant of ACT [6], which we refer to as Geometric Compliant ACT (G-CompACT). This policy handles detailed motion and contact interaction using force-torque feedback and RGB inputs from wrist-mounted cameras. A key design feature of G-CompACT is that it only relies on local information—specifically, the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs. The output of G-CompACT is the desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control. We emphasize that our contribution is on the structural framework rather than the specific choice of algorithms. For example, one could replace the Diff-EDF with other equivariant methods, such as ET-SEED [15], or replace the ACT with the DP or their variants, e.g., Diffusion Transformer (DiT) [16]. The main contributions of this paper are as follows:
@@
    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability in real robot experiments involving tight insertions.
@@
    We identify three key principles for spatially generalizable contact-rich manipulation: Compliance, Localized Policy (Left invariance), and Induced Equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
@@ II Related Works
Equivariant Methods Equivariant methods aim to leverage the inherent symmetries of robot tasks, typically modeled as group transformations such as $SE(3)$, to improve the sample efficiency of imitation learning. Early approaches, such as Neural Descriptor Fields (NDFs) [22] and Transporter Networks [3], incorporated equivariance to $SE(3)$ and $SO(2)$ transformations, utilizing scene point clouds and top-down views, respectively. Equivariant Descriptor Fields (EDFs) [23] reformulated the NDF approach into a fully end-to-end learning method with $SE(3)$ bi-equivariance property, which is further improved by Diffusion-EDF (Diff-EDF) [14] to improve training and inference performance. Other equivariant models that extract desired keyframes from point clouds include [24]. 3D Equivariant extensions of DP [25, 26, 15] and Flow-Matching [27, 28] policies have been explored in recent literature and combine $SE(3)$ equivariant observation encoders with appropriately modified diffusion/flow-matching processes. Compared to recent works of [29, 30], we generalize the table-top manipulation to contact-rich tasks while achieving full $SE(3)$ equivariance from vision to force/control level, bridging vision, control, and force interaction under a unified framework.
Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks. The proposed EquiContact consists of G-CompACT and Diffusion-EDF. The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen scenarios during evaluation.
@@ III Problem Definition
In this paper, we aim to identify the key structural components of a policy capable of executing spatially generalizable, contact-rich manipulation tasks. As mentioned earlier, our proposed building blocks to achieve $SE(3)$ vision-to-force equivariant policy are: 1. Compliant Control, 2. Localized Policy, and 3. Induced Equivariance - which we will experimentally validate through specific data collection and evaluation setups, as detailed below.
III-A Problem Setup
Our primary focus is to provide a learning-based solution for a peg-in-hole task,a classic representative of contact-rich, force-based robotic manipulation and assembly tasks, using vision. Unlike prior work such as [11], which assumes that the desired hole pose is known and that the robot starts with a pre-grasped peg, we consider a more general scenario. The robot first needs to pick up the peg and then insert it into the hole, as shown in Fig. 2, relying solely on vision and proprioception. We assume only that the peg is upright in the scene. Following the reasoning that every manipulation task can be regarded as a sequence of pick-and-place, we formulated peg-in-hole insertion as a compliant pick-and-place task. However, as highlighted in the earlier discussion, naively using the pick-and-place approach is insufficient because the high-level vision planner often lacks the accuracy required for precise and contact-rich tasks, which is $1\mathrm{mm}$ in our case (see Fig. 2).
To accommodate this, we propose a low-level policy that:
  • provides real-time visual feedback to correct errors from the coarse high-level planner,
@@ III-A Problem Setup
Assuming that a high-level vision planner that returns approximate desired end-effector pick-and-place poses is readily available, we focus on developing equivariant compliant placing, i.e., an insertion policy via imitation learning, and regard picking as a spatial case of placing.
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location. This enables the policy to perform nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. $2$ external cameras with calibrated extrinsics and $2$ wrists cameras are installed. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
@@ IV Solution Approach
We now describe our proposed EquiContact policy, which integrates a high-level vision-based planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant control Action Chunking Transformer (G-CompACT). G-CompACT itself consists of a behavior cloning module built on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC). As discussed in Section III, because of space constraints, we focus on the insertion task in this paper, which will later be extended to include the picking task to complete the full pipeline implementation. We note that the Diff-EDF and G-CompACT are trained separately but are combined as a whole pipeline during inference time, as shown in Fig. 1.
@@ IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that:
$g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
@@
To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background. This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene.
While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and have high robustness to left $SE(3)$ transformations.
@@
While the original Diff-EDF supports right-equivariance to handle transformations between the gripper and the grasped object, we did not utilize this feature. In practice, we observed that arbitrary peg transformations relative to the gripper introduced additional challenges. First, grasping the peg based on an arbitrary transformation leads to imprecise grasps, resulting in object slippage during contact. Second, the gripper-object rotation must be estimated continuously in real-time for G-CompACT to work under slippage. Instead, we enforced a consistent grasp orientation and relied only on the left-equivariance property. We refer the readers to [14, 10] for full details of Diff-EDF and to Supplementary Material  SI for implementation and heuristics. We emphasize again that Diff-EDF can be replaced by any point cloud-based method that outputs an approximate goal pose for (compliant) placing.
@@ IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
IV-B1 Geometric Admittance Control (GAC)
We implement the geometric impedance control (GIC) proposed in [12, 13] in the geometric admittance control (GAC) setup [11]. We first note that we follow the notation used in [11]. Let the end-effector pose be denoted as $g\in SE(3)$ in a homogeneous matrix representation, or simply $g=(p,R)$, where $p\in\mathbb{R}^{3}$ is a position of the end-effector and $R\in SO(3)$ is a rotation matrix of the end-effector. Given the desired end-effector pose $g_{d}=(p_{d},R_{d})$, the desired end-effector dynamics for the GAC setup is given as follows:
$M\dot{V}^{b}+K_{d}V^{b}+f_{{}_{G}}=F_{e},$ (3)
where $M\in\mathbb{R}^{6\times 6}$ is symmetric positive definite desired inertia matrix, $K_{d}\in\mathbb{R}^{6\times 6}$ symmetric positive definite damping matrix, $F_{e}\in\mathbb{R}^{6}$ is external wrench applied to the end-effector in end-effector body frame and $V^{b}\in\mathbb{R}^{6}$ is a body-frame end-effector velocity. Further, $f_{{}_{G}}=f_{{}_{G}}(g,g_{d},K_{p},K_{R})\in\mathbb{R}^{6}$ is a elastic wrench given by:
$\begin{split}f_{{}_{G}}=\begin{bmatrix}f_{p}\\ f_{R}\end{bmatrix}=\begin{bmatrix}R^{T}R_{d}K_{p}R_{d}^{T}(p-p_{d})\\ (K_{R}R_{d}^{T}R-R^{T}R_{d}K_{R})^{\vee}\end{bmatrix},\end{split}$ (4)
where $\hat{f}_{{}_{G}}\in se^{*}(3)$, with $K_{p},K_{R}\in\mathbb{R}^{3\times 3}$ symmetric positive stiffness matrices for the translational and rotational dynamics, respectively, and $(\cdot)^{\vee}$ denotes the vee-map. By a slight abuse of notation we will frequently denote as $f_{{}_{G}}\in se^{*}(3)$ to emphasize that it is an elastic wrench on $SE(3)$, since $f_{{}_{G}}$ is isomorphic to $\hat{f}_{{}_{G}}$ [13].
Given the desired dynamics (3), the desired end-effector pose command $\tilde{g}_{d}(k)$ provided to the end-effector controller is calculated in discrete time as
$V_{d}^{b}(k)\!$ $=\!V^{b}(k)\!+\!T_{s}\cdot M^{-1}(F_{e}(k)\!-\!f_{{}_{G}}(k)\!-\!K_{d}V^{b}(k)),$
$\tilde{g}_{d}(k)\!$ $=\!g(k)\cdot\exp{(\hat{V}_{d}^{b}(k)\cdot T_{s})},$ (5)
where $T_{s}$ is a sampling time and $\hat{(\cdot)}$ denotes a hat-map. For the details of the GIC/GAC, we refer to [12, 11]. The admittance control loop is implemented at a $200\mathrm{Hz}$ rate using ROS2. We refer to Supplementary Material SI for detailed implementation tips.
IV-B2 G-CompACT
Action Chunking with Transformer (ACT) is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks, originally introduced in [6]. While the original ACT formulation encodes robot proprioception using joint positions, this representation lacks spatial equivariance, limiting generalization. Instead, we adopt a proprioceptive encoding based on the end-effector pose, following the approach in [17].
In our previous work [11], we proposed a recipe for spatially equivariant policies, which involves a left invariance property and policy representation in the end-effector body frame. Following this principle and recent developments on relative action representations (similar to [29]), we structure our G-CompACT policy to take observations composed of: (1) Geometrically Consistent Error Vector (GCEV) proposed in [11], (2) FT sensor in the end-effector frame to capture contact behaviors, and (3) RGB images from wrist cameras. The model returns as action sequences: (1) relative pose from the current end-effector frame, and (2) admittance gains.
The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as

### v2

@@ Abstract
This paper presents a framework for learning vision-based robotic policies for contact-rich manipulation tasks that generalize spatially across task configurations. We focus on achieving robust spatial generalization of the policy for the peg-in-hole (PiH) task trained from a small number of demonstrations. We propose EquiContact, a hierarchical policy composed of a high-level vision planner (Diffusion Equivariant Descriptor Field, Diff-EDF) and a novel low-level compliant visuomotor policy (Geometric Compliant Action Chunking Transformers, G-CompACT). G-CompACT operates using only localized observations (geometrically consistent error vectors (GCEV), force-torque readings, and wrist-mounted RGB images) and produces actions defined in the end-effector frame. Through these design choices, we show that the entire EquiContact pipeline is $SE(3)$-equivariant, from perception to force control. We also outline three key components for spatially generalizable contact-rich policies: compliance, localized policies, and induced equivariance. Real-world experiments on PiH, screwing, and surface wiping tasks demonstrate a near-perfect success rate and robust generalization to unseen spatial configurations, validating the proposed framework and principles. The experimental videos can be found on the project website: https://sites.google.com/berkeley.edu/equicontact
@@ I Introduction
However, such policies often lack spatial generalizability and therefore require a large amount of data to learn robust behaviors. As described in [9], both action chunking transformers (ACT) [6] and diffusion policy (DP) [5] are evaluated only within the limited spatial variations. Furthermore, both methods exhibit near-linear performance growth as the demonstration dataset size increases, suggesting that the trained policies do not inherently generalize well to new spatial configurations, but rather tend to interpolate between seen demonstrations.
@@
In [11], a $SE(3)$-equivariant gain-scheduling policy using geometric impedance control (GIC) [12, 13] was proposed to solve peg-in-hole (PiH) problems. Inspired by the view that many manipulation tasks can be framed as pick-and-place problems [4], we modeled PiH as a compliant pick-and-place task, where final peg poses are provided by vision-based $SE(3)$-equivariant models such as Diffusion-EDF (Diff-EDF) [14]. Since both the high-level planner and low-level variable impedance controller are equivariant, they can be combined to form a vision-to-force equivariant policy. However, in practice, Diff-EDF’s placement accuracy proved insufficient for tight insertion tasks, which require sub-millimeter precision (details in Appendix -A). This revealed a key limitation: high-level vision planners may capture global structures but struggle with precision and contact-sensitive execution. From this insight, we introduce an intermediate layer between the planner and the low-level controller, which provides real-time visual feedback to correct the residual errors of the high-level planner.
In this paper, we propose EquiContact, a hierarchical $SE(3)$ vision-to-force equivariant policy for spatially generalizable, contact-rich tasks. It consists of two main components: a high-level planner using Diffusion Equivariant Descriptor Fields (Diff-EDF) [14], which estimates a local reference frame from point clouds, and a low-level compliant visuomotor policy based on Action Chunking Transformer (ACT) [6], which we refer to as Geometric Compliant ACT (G-CompACT). A key design feature of G-CompACT is that it only relies on local information: the force-torque signal in the end-effector frame, a geometrically consistent error vector (GCEV) [11], and wrist camera inputs. The output of G-CompACT is the relative desired pose and admittance gains, which are then sent to the geometric admittance controller (GAC) module to execute compliant control. Our contribution lies in the framework design, not in specific model choices; for example, Diff-EDF could be replaced by ET-SEED [15], or ACT by other visuomotor policies.
The main contributions of this paper are as follows:
@@
    We identify three key principles for spatially generalizable contact-rich manipulation: left-invariant compliant control action (via GAC [11]), localized policy (left invariance), and induced equivariance. These enable $SE(3)$-equivariant behavior without requiring explicitly equivariant neural networks [10].
@@
    Under these principles, we present the necessary conditions for $SE(3)$ vision-to-force equivariant policy, and mathematically prove the equivariance property of EquiContact.
  4. 4.
    We demonstrate that EquiContact achieves near-perfect success rates and spatial generalizability when these conditions are met in real robot experiments involving peg-in-hole, screwing, and surface wiping tasks.
@@ II Related Works
Equivariant Methods Earlier equivariant approaches attempted to handle manipulation tasks as an extension of pick-and-place tasks, by leveraging $SE(3)$ equivariance from point clouds [22, 15, 14] or $SO(2)$ equivariance [3] from top-down views. Equivariant approaches have been extended to visuomotor policies, such as DP or flow matching, [23, 24], using point clouds. [25, 26] proposed $SO(2)$ equivariant visuomotor policies using 2D images, not fully considering $SE(3)$. In contrast, our approach induces full $SE(3)$ equivariance from vision to control force without relying on explicitly equivariant neural networks, but using structured observations and actions via geometrical canonicalization. Furthermore, by integrating with $SE(3)$ equivariant control, we generalize beyond table-top settings to contact-rich manipulation.
Manipulation in Object Frame Our induced $SE(3)$ equivariant approach relies on representing the visuomotor policy in the end-effector frame. While recent works [27, 28, 29] define policies in the target (object) frame, policy representation in the end-effector frame offers improved fidelity and robustness. This is because the estimated object frame can be noisy, and the end-effector frame is reliably obtained via forward kinematics. Importantly, compared to [27, 28], we explicitly link the choice of reference frame to the equivariance property, and unlike [29], which only handles translational transformations, our method can deal with full $SE(3)$ transformations of the reference frame.
Figure 1: We propose an EquiContact, a hierarchical, provably $SE(3)$ vision-to-force equivariant policy for spatially generalizable contact-rich tasks. The proposed EquiContact consists of G-CompACT (Section. IV-B) and Diffusion-EDF (Section. IV-A). The G-Compact plays a localized policy over the reference frame provided by the Diffusion-EDF, making our framework generalizable to unseen task transformations during evaluation. The overall EquiContact algorithm is summarized in Algorithm 1.
@@ III Problem Definition
In this paper, we aim to identify the key structural components required for learning policies that generalize spatially in contact-rich manipulation tasks. We will first focus on the peg-in-hole (PiH) problem as a representative force-based assembly task and validate the feasibility of the proposed approach to other contact-rich tasks later. Our proposed framework achieves $SE(3)$ vision-to-force equivariance through three essential design principles: (1) left-invariant compliant control, (2) localized policy, and (3) induced equivariance. These principles are validated through specific data collection and evaluation setups, as detailed below.
Unlike prior work [11] that assumes a known hole pose and a pre-grasped peg, we consider a more general setup: the robot must first grasp the peg and then perform insertion using only vision and proprioception, as illustrated in Fig. 2. We assume only that the peg is upright in the scene. Inspired by the idea that many manipulation tasks can be viewed as sequences of pick-and-place actions, we formulate insertion as a compliant pick-and-place task. However, high-level vision planners often lack the precision needed for tight-tolerance tasks like PiH, which in our case requires $1\mathrm{mm}$ accuracy.
To accommodate this, we propose a low-level compliant policy that:
  • provides real-time visual feedback to refine the coarse high-level command,
@@
We assume that a high-level vision planner (e.g., Diff-EDF) can generate approximate pick-and-place poses. Our focus is on developing an equivariant compliant placing policy, i.e., insertion policy, using imitation learning. We further assume the peg is approximately aligned with the gripper during placement, since arbitrary peg poses introduce two challenges: (1) imprecise grasps can lead to slippage during contact, and (2) compensating for slippage requires continuous estimation of the gripper-to-peg transformation, which is difficult to achieve reliably in real time.
To train this policy, we collect expert demonstrations of insertions on a fixed-platform setting with a known hole location, where its objective is to train a policy that performs nearly perfectly in the trained scenario. We then evaluate benchmark and proposed methods—trained solely on these limited demonstrations—across arbitrarily translated and rotated test scenarios, thereby isolating and testing individual components of our spatially generalizable contact-rich policy.
Figure 2: (Left) Overview of the workspace for the peg-in-hole assembly task is presented. 22 external cameras with calibrated extrinsics and 22 wrists cameras are installed. (Right-Top) Peg and hole assembly with $1\mathrm{mm}$ of clearance. (Right-Bottom) Hole part with flat and tilted ($30^{\circ}$) platforms.
@@ IV Solution Approach
We now describe our proposed EquiContact policy, which integrates a high-level vision planner—Diffusion Equivariant Descriptor Field (Diff-EDF)—with a low-level compliant visuomotor policy—Geometric Compliant ACT (G-CompACT). G-CompACT consists of a behavior cloning module based on a transformer-based CVAE architecture, followed by a geometric admittance controller (GAC). As discussed in Section III, we focus on the insertion task, with extension to picking addressed later in the paper. The Diff-EDF and G-CompACT are trained separately and combined at inference, as shown in Fig. 1.
Conceptually, EquiContact follows a simple yet powerful principle: “anchoring a localized policy on a globally estimated reference frame.” In our framework, G-CompACT serves as the localized low-level policy, operating solely on observations defined in the current end-effector frame; wrist camera RGB images, force-torque (FT) signals, and a geometrically consistent error vector (GCEV) $e_{{}_{G}}$. It outputs the desired relative pose and admittance gains, also in the end-effector frame. As both inputs and outputs are defined locally, G-CompACT functions as a fully localized policy. The high-level planner, Diff-EDF, estimates the pose of the target (e.g., hole) in the world frame. At the inference, the robot moves near the estimated reference frame, and G-CompACT is activated to perform compliant motion using only local feedback. Because the low-level policy does not depend on absolute global inputs, it can transfer robustly to unseen spatial configurations when the estimated reference frame is provided. This divide-and-conquer design provides a general framework for enhancing both spatial generalization and policy interpretability for contact-rich, and more broadly, general manipulation tasks.
In the remainder of this section, we formalize this spatial generalization property as $SE(3)$ equivariance and show how EquiContact satisfies this by construction through a rigorous mathematical derivation. We note that $\theta$, $\theta_{i}$, and $\phi$ with $i=1,2,3$ denote a learnable parameter of the neural network, without loss of generality. We also note that we follow the notation and definition conventions of [11, 30].
@@ IV-A Diffusion-Equivariant Descriptor Field (Diff-EDF)
where the neural network $f_{\theta_{1}}$ implements the mapping. Due to the left-equivariant property, it follows that: $g_{l}g_{{}_{EDF}}=f_{\theta_{1}}(g_{l}\circ\mathcal{O}^{scene},\mathcal{O}^{grasp}).$
@@
To meet this assumption, we randomize the pose of the hole-platform assembly during training, both translationally and rotationally, with visual distractors in the background. This encourages the model to focus on $\mathcal{O}^{ref}$, not the rest of the objects in the scene. While this property is not strictly enforced (due to the learned nature of $f_{\theta_{1}}$), Diff-EDF utilizes equivariant backbone architectures and localized attention mechanisms [31], enabling it to generalize from as few as $\sim 10$ demonstrations and has high robustness to left $SE(3)$ transformations.
@@
Note that arbitrary peg transformations relative to the gripper can be handled via the right-equivariance feature of the Diff-EDF; however, due to the complications mentioned in Section. III, we enforced a consistent grasp and only utilized the left-invariance property. We refer the readers to [14, 10] for full details of Diff-EDF.
@@ IV-B Geometric Compliant control Action Chunking with Transformers (G-CompACT)
IV-B1 G-CompACT
G-CompACT is based on the Action Chunking with Transformer (ACT), which is a CVAE-based generative model designed for imitation learning in robotic manipulation tasks [6]. To make G-CompACT spatially equivariant, we follow the principles proposed in [11]: left-invariant policy and policy representation in the end-effector body frame. To achieve this, we structure the observations and actions of the G-CompACT as described below.
First, the observations are given by, (1) Geometrically Consistent Error Vector (GCEV) $e_{{}_{G}}$ proposed in [11], (2) FT sensor in the end-effector frame $F_{e}$ to capture contact behaviors, and (3) RGB images $I_{w}=\{I_{w,1},I_{w,2}\}$ from wrist cameras (see Fig. 3). For actions, we choose (1) relative pose from the current end-effector frame $g_{rel}$, and (2) admittance gains for Geometric Admittance Control (GAC) $(K_{p},K_{R})$. The details of GAC and the definition of gains will be provided later in the Section.
Formally, let the G-CompACT method be $\pi_{\theta_{2}}$ such that
$\begin{split}a(k)&=\pi_{\theta_{2}}(o(k)),\quad\text{where}\\ a(k)&\triangleq(g_{rel},K_{p},K_{R})(k),\quad o(k)\triangleq(e_{{}_{G}},F_{e},I_{w})(k),\end{split}$ (3)
where $a(k)$ denotes the actions, and $o(k)$ denotes the observation at time step $k$. In fact, our G-CompACT also outputs the actions of chunk size $N$ following the standard action-chunking strategy. However, for the compact presentation, we will only consider the single-step action after undergoing proper processing, such as a temporal ensemble. The GCEV $e_{{}_{G}}(g,g_{{}_{EDF}})$ is defined as